Write an equation in standard form with center at the origin and with the given characteristics. Ellipse height 15 units , width 8 units
Answer by teishad(7)

(x^2)/4^2 + (y^2)/7.5^2 = 1
that would be your equation but i would check and make sure that the number for up and down is really 15 because they usually dont have them as in between numbers. make sure its not 14 or 16
if its 14 it would be
(x^2)/4^2 + (y^2)/7^2 = 1
and if its 16 it would be
(x^2)/4^2 + (y^2)/8^2 = 1
feel free to email me if you have any more questions, i love these!!! =)
tdiers@frontier.com